Wardell Company purchased a mainframe on January 1, 2016, at a cost of $40,000. The computer was depreciated using the straight-line method over an estimated five-year life with an estimated residual value of $4,000. On January 1, 2018, the estimate of useful life was changed to a total of 10 years, and the estimate of residual value was changed to $400.

Required:
1. Prepare the year-end journal entry for depreciation in 2018. No depreciation was recorded during the year.
2. Prepare the year-end journal entry for depreciation in 2018. Assume that the company uses the sum-of-the-years' -digits method instead of the straight-line method.
Answer #1

Given:

Cost of mainframe as on January 1 2016 = $40,000

Salvage value = $4,000

Useful life = 5 years

Depreciation is calculated using straight line method as shown below:

Depreciation = Cost-Salvage value Useful life 40,000-4,000 = $7,200

Depreciation for 2016 and 2017 = 7,200 * 2 = $14,400

Book value as on January 1 2018 = Cost – Depreciation for 2017 and 2018

= 40,000 – 14,400

= $25,600

On January 1 2018, it was estimated that useful life of mainframe was 10 years and salvage value was $400.

Amount of depreciation in 2018 will change accordingly. Useful life will be 8 years (10 – 2) as 2 years have already been accounted for.

Journal entry to record depreciation for 2018 :

Date

Particulars

Debit

Credit

31 December 2018

Depreciation expense

$3,150

Accumulated depreciation

$3,150

(To record depreciation expense)

2. Calculation of depreciation using sum of digits method:

Depreciation - Cost-Salvage valuex Dedining years Sum of years

Sum of years = 5 + 4 + 3 + 2 + 1 = 15

Substitute the values in the above formula:

Depreciation in 2016 = (40,000 – 4,000) * 5 / 15

= $12,000

Depreciation in 2017 = = (40,000 – 4,000) * 4 / 15

= $9,600

Book value on January 1 2018 = Cost – depreciation for 2016 and 2017

= 40,000 – 12,000 – 9,600

= $18,400

Useful life from 2018 onward is 8 years (10 – 2)

Sum of years = 8 + 7 + 6 + 5 + 4 + 3 + 2 + 1 = 36

Salvage value is $400

Depreciation in 2018 = (18,400 – 400) * 8/36

= $4,000

Journal entry to record depreciation for 2018 :

Date

Particulars

Debit

Credit

31 December 2018

Depreciation expense

$4,000

Accumulated depreciation

$4,000

(To record depreciation expense)